A 500-W heating coil designed to operate from 110 V is made of Nichrone wire 0.0500 mm in diameter a) assuming the resistivity of the Nichrone remains constant at its 20.0 degrees celcius , find the length of wire used. b) What if? Now consider the variation of resistivity with temperature . what power is delivered to the coil of part a) when is is warmed to 1200 degrees celcius?
1
Expert's answer
2020-04-08T10:41:57-0400

a)


"R=\frac{U^2}{P}=\rho\frac{l}{A}=\rho\frac{4l}{\pi d^2}"

"\frac{110^2}{500}=\frac{4(100\cdot10^{-8})}{\pi (5\cdot10^{-5})^2}l"

"l=0.0475\ m"

b)


"P'=\frac{P}{1+\alpha(T-T_0)}"

"P'=\frac{500}{1+0.0004(1200-20)}=340\ W"